solve the differential equation by reduction of order:

= x2y + 2xy 6y= 0 given that yn=x?

could use some help solving this
Answer #1

11 = x2 - x2y 1 + 2xy _by=0 - Sol put y = Y, Y = Y= vx2 - y= x²vt aur - y zavt x²/ + zu vt av y = x2v1 + 40 +2V - SubN = Goto =) SdV=Glatban v=9s +2 V=-9 85+ (2 so y = vy so goo= - 1 x ²4 C₂22 orf 41x)= - - 73+ (2x ?